# A new long gamma-ray burst formation pathway at solar metallicity

### v1

@@ 3.1.3 Metallicity dependence
At $Z=0.45\leavevmode\nobreak\ \mathrm{Z}_{\odot}$, the region with sufficient angular momentum at collapse in the stable reverse-mass-transfer channel shrinks, as shown in Figure 4. Weaker main-sequence winds and partial stripping of the primary star during the first mass transfer phase (Götberg et al., 2017) lead to the formation of a larger helium core and a small hydrogen envelope at the reverse mass transfer. Similar to large-period binaries at $\mathrm{Z}_{\odot}$, this results in a larger primary radius and less angular momentum gain during the reverse mass transfer. Additionally, the larger helium star loses more mass and angular momentum before collapse, preventing accretion disk formation. At the same time, the islands of unstable mass transfer from $L_{2}$ overflow and the maximum $\dot{M}$ are expanding.
Moving to $Z=0.2\leavevmode\nobreak\ \mathrm{Z}_{\odot}$, two additional effects are starting to show: short-period systems no longer experience reverse mass transfer, while wider-period systems are now experiencing a contact phase, which becomes dominant towards even lower metallicities, as the $Z=10^{-3}\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ panel in Figure 4 shows. In wider binaries, the mass transfer no longer sufficiently strips the primary star for it to reach the Wolf-Rayet wind regime, instead, it finds a thermal equilibrium close to or at its Roche lobe (Klencki et al., 2022; Ercolino et al., 2024). As the companion evolves off its main sequence and also fills its Roche lobe, the system goes into contact, either because the primary is already filling its Roche lobe or because its equilibrium state is near the Roche lobe, such that the radial expansion from mass accretion causes the primary to fill its Roche lobe. On the other hand, short-period binaries no longer experience reverse mass transfer, because the rejuvenated companion remains compact and hydrogen-burning until the primary reaches carbon exhaustion. Due to lower stellar wind mass loss, these accretors retain the gained angular momentum from the initial mass transfer phase. At lower metallicities, the radial expansion of the stars is already less compared to solar metallicity due to the lower opacity, and the increased rotational mixing further limits the radial expansion, which restricts a reverse-mass-transfer phase. The combination of these effects creates a LGRB formation channel that becomes more efficient as metallicity increases, while completely disappearing at metallicities below $0.2\leavevmode\nobreak\ \mathrm{Z}_{\odot}$.
@@ 3.2 Stable Reverse Mass Transfer LGRB Rate
Given that most observed LGRBs occur in low-metallicity environments, this indicates that some effect must be limiting a large fraction of our stable reverse mass-transfer model progenitors from producing a LGRB. Theoretically one would expect that such an effect follows from the required energy for the jet to propagate through the stellar envelope and successfully break out of it (e.g. Matzner, 2003; Bromberg et al., 2011; Salafia et al., 2020). The binding energy of the stellar structure at core carbon exhaustion is less than the values of $E_{\mathrm{BZ}}/\eta_{\phi}$, when $\eta_{\phi}=1$. While the actual energy breakout in general depends on the properties of the stellar envelope and also on those of the jet at launch. Here, for simplicity, we model both as a universal threshold on the jet energy. Because we left the efficiency factor $\eta_{\phi}$ unspecified (see Section 2), we set an energy limit for a successful LGRB of $10^{51}\leavevmode\nobreak\ \mathrm{erg}$ based on the typical breakout energy from jet propagation simulation (for example, see Harrison et al., 2018). In what follows, we explore five values for the efficiency, $\eta_{\phi}$, namely $1$, $0.1$, $10^{-2}$, $10^{-3}$, and $5\times 10^{-4}$.
The total rate of disk formation through the stable reverse-mass-transfer channel is comparable to the local rate density of observed LGRBs (Ghirlanda & Salvaterra, 2022). Given that the inferred fraction of high-metallicity host galaxies is 10%-20%, this requires an efficiency that removes 80%-90% of events. The most restrictive efficiency with $\eta_{\phi}=5\times 10^{-4}$ results in a volumetric rate of ${\simeq}1.83$ Gpc-3 yr-1 at $z=0$; with a total of ${\simeq}13.3$ Gpc-3 yr-1 events below $z<2.5$. With $\eta_{\phi}=10^{-3}$, the reverse-mass-transfer channel rate is around 18% of the latest estimate of the observed cosmic LGRB rate. This is remarkably close to the observational fraction of solar and super-solar host galaxies of 10-20 per cent. However, because the normalisation of the observed rate is uncertain due to selection effects, such as the intrinsic energy of the jet, one should be careful to interpret a comparison of our theoretical rates against these observations. These cannot be constrained without considering all formation channels concurrently, which is beyond the scope of this work. The intrinsic rate density evolution with redshift for different choices in the efficiency is shown in Figure 5. We report their rate densities at $z=0$ and averaged within $z=2.5$ in Table 1.
@@
Figure 5: The redshift evolution of the stable reverse-mass-transfer LGRB event rate density for the IllustrisTNG metallicity and star formation history. The total potential population is indicated in black, while the dashed red line shows the intrinsic LGRB rate from Ghirlanda & Salvaterra (2022) with the best-fit parameters based on observations. The five green shaded lines indicate the rate density for an energy limit of $10^{51}\leavevmode\nobreak\ \mathrm{erg}$ and different $\eta_{\phi}$.
@@ 4 Discussion
While LGRBs have been found in high-metallicity environments, no collapsar formation channel for such events has previously been proposed. We have introduced a new formation channel capable of forming sufficient LGRBs to match observational or theoretical fractional rates in high-metallicity environments. However, much uncertainty remains in this fraction, as well as in the absolute rate of the total LGRBs and this channel. Since the main contributors to this uncertainty are the observational biases and hidden LGRB fraction, which would require a population study that includes all possible LGRB formation pathways in order to be explored, we instead focus on the physics which might impact the presence and occurrence of the reverse-mass-transfer LGRB formation channel.
@@ 4.1 Implications of energy limit
We have implemented a universal energy limit as a proxy for successful jet breakout with different efficiencies, which is equivalent to different energy limits with a constant efficiency, i.e. a $10^{51}\leavevmode\nobreak\ \mathrm{erg}$ limit with a $\eta_{\phi}=10^{-3}$ select the same progenitors as a $10^{54}\leavevmode\nobreak\ \mathrm{erg}$ limit with a $\eta_{\phi}=1$. For the stable-reverse-mass-transfer channel to produce 10-20% of the local LGRB rate, a low disk-to-jet-energy conversion or a high jet-breakout-energy threshold is required. This implies that successful jets are a minority, while choked jets occur an order of magnitude more frequently in the proposed channel. This fraction aligns with Bromberg et al. (2012) and from observed low-luminosity GRBs the inferred intrinsic rate can be up to an order of magnitude higher than their high-luminosity counterpart (Liang et al., 2007; Guetta & Valle, 2007; Virgili et al., 2009). It suggests the existence of a large population of relativistic supernovae with potentially low-luminosity gamma-ray emission caused by these choked jets (see, for example, Soderberg et al., 2006; Wang et al., 2007; Nakar & Sari, 2012).
@@ 4.2 Type Ic Broad-line and potential LGRB companions
LGRBs and SN Ic-BL have been associated with one another, but this relationship is not universal; not all Type Ic-BL are accompanied by a LGRB and vice versa (Modjaz et al., 2016; Japelj et al., 2018). Type Ic-BL without a LGRBs have distinct spectral features from those with a LGRB (Modjaz et al., 2016), and could indicate a failed jet breakout (see Corsi & Lazzati, 2021, and references therein). In general, Type Ic-BL are associated more with super-solar host galaxies compared to LGRBs (Japelj et al., 2018), which aligns with the positive metallicity dependence of the stable reverse-mass-transfer LGRB channel. Additionally, if this channel produces the observed high-metallicity population of LGRBs, it produces an order of magnitude more choked jets compared to successful jets (see Section 4.1). Although the observed relative rate of SN Ic-BL to core-collapse SNe is poorly constrained, based on a single SN Ic-BL observation in the volume-limited Lick Observatory Supernova Search, SN Ic-BL make up 4% of stripped-envelope supernovae (Shivvers et al., 2017). Combined with the stripped-envelope supernova rate from Frohmaier et al. (2021), this results in a rough SN Ic-BL rate estimate of ${\sim}300$ Gpc-3 yr-1. Assuming that all disk forming systems produce a jet and $\eta_{\phi}=10^{-3}$, ${\simeq}86$ choked jets occur per Gpc3 per year at $z=0$ from the stable reverse-mass-transfer channel. This is comparable to the observationally inferred rate and could imply that either more dominant LGRB formation channels must have a smaller ratio of choked-to-successful jets to not overproduce the SN Ic-BL or not all failed LGRB accretion disk systems produce a SN Ic-BL.
@@ 4.3 Jet Physics
The relationship between the progenitor star and jet energy remains highly uncertain and general relativistic, radiation, magnetohydrodynamic simulations have shown that the efficiencies depend on the state of the disk (Tchekhovskoy et al., 2010). A popular approach is the magnetically arrested disk (MAD) state because it produces powerful jets even for slowly rotating BHs and causes rapidly spinning BHs to spin down (Jacquemin-Ide et al., 2024). In this state, the mass accretion rate and spin-down timescale determine the jet power. In such disks $\eta_{\phi}=1$, while $\eta_{a}$ can be fitted following Lowell et al. (2024). This can result in efficiencies above 1 due to the extraction of angular momentum from the BH and results in high-energy jets. These even occur for slowly spinning BHs as long as the progenitor has sufficient angular momentum to form an accretion disk (Fryer et al., 2024). However, recent simulations of collapsars show that the accretion rate remains sufficiently high that neutrino cooling keeps the disk thin and the MAD state is not reached when assuming a proto-neutron star is the origin of the BH magnetic field (Gottlieb et al., 2024).
@@
Besides the potential energy budget to form a jet, an essential component in LGRB formation in the collapsar scenario is the jet breakout through the stellar structure. The jet breakout depends on the injected energy and active duration of the central engine, and the density profile of the stellar structure. While the internal stellar density profile at core-carbon exhaustion is available, we do not evolve the final evolutionary phases due to the required computational time, and, thus, do not have a representative profile at collapse available. We verify that the free-fall timescales of the core-carbon exhausted models are similar to observed $T_{90}$ values (See Appendix B). Furthermore, the injected energy depends on several poorly constrained parameters, such as the injection opening angle, the disk mass to jet energy conversion, and the total jet energy emitted in the gamma-ray. A full end-to-end pipeline with computationally intensive hydrodynamical simulations is required to accurately model the link between the progenitor model and the jet properties from the collapsar scenario (Gottlieb et al., 2024).
@@ 4.4 Star Formation History
Although the total star formation rate density in the local Universe is relatively well constrained, the metallicity evolution and spread are less understood. It is the high-metallicity-specific star formation rate that gives rise to uncertainty in the shape of the stable reverse-mass-transfer LGRBs and in the weighting of each model over redshift. The former can be seen in Figure 7, where we compare the LGRB rates over redshift from the IllustrisTNG simulation with the two empirical star formation history prescriptions described in Section 2. The metallicity distribution redistributes the weight between populations of different metallicities but does not alter the shape of the potential LGRB energy distribution per metallicity (see Appendix A). The ln-normal metallicity distribution from Neijssel et al. (2019) produces a very peaked metallicity-specific star formation rate, reducing the contribution from the reverse-mass-transfer LGRB channel at high redshift to the LGRB rate, while the higher metallicity and wider distribution from Bavera et al. (2020) allows for more near-solar metallicity star formation at higher redshift.
@@
A stronger effect in the local Universe comes from the energy cuts. Due to the reweighing of each LGRB progenitor, the cumulative potential energy distribution is significantly different per star formation history, thus, the energy cut can suddenly remove more events. For example, the most restrictive cut in Figure 6 reduces the stable reverse-mass-transfer LGRB rate to 1.83 LGRB Gpc-3 yr-1, while the Neijssel et al. (2019) prescription in Figure 12 reduces it to 0.12 LGRB Gpc-3 yr-1. This is due to the energy distribution at each metallicity, which each SFH rebalances, as can be seen in Appendix A. As a result, the choice in mean metallicity and metallicity spread has a major impact on the final rate and the limit of the jet energy.
@@ 4.5 Additional Caveats
Although the majority of LGRB progenitors from the stable reverse-mass-transfer channel are well into the BH formation regime, its parameter space at low masses is limited by neutron star formation, as shown in the right grid slice in Figure 3. Allowing for more BH formation would increase the contribution of this formation channel. However, the additional parameter space is limited, because the timescale difference between the binary components is too large at low masses for reverse mass transfer to occur. Additionally, the Patton & Sukhbold (2020) prescription uses the average central carbon abundance and carbon-oxygen core size at helium depletion to determine the explodability. Late-stage binary interactions, such as mass transfer after core helium ignition, can alter the stellar structure sufficiently to boost explodability (Laplace et al., 2021). Despite that, the POSYDON MESA models are evolved until carbon depletion, due to the Patton & Sukhbold (2020) prescription late stage mass transfer effects are not considered.
Switching the remnant mass prescription to the rapid or delayed prescriptions from Fryer et al. (2012) should give an indicator of how different prescriptions for estimating the remnant masses affect our results. These prescriptions use the core properties at carbon depletion to determine BH formation through fallback onto a proto-compact object. Figure 8 shows minimal differences between their metallicity bias functions, which propagates to nearly the same cosmic rates. Slightly less stable reverse-mass-transfer LGRBs occur with the Fryer et al. (2012) prescriptions. The star formation history, metallicity distributions, and energy cut have a significantly stronger impact on this rate than what stars collapse as BHs.

### v2

@@ 3.1.3 Metallicity dependence
At $Z=0.45\leavevmode\nobreak\ \mathrm{Z}_{\odot}$, the region with sufficient angular momentum at collapse in the stable reverse-mass-transfer channel shrinks, as shown in Figure 4. Weaker main-sequence winds and partial stripping of the primary star during the first mass transfer phase (Götberg et al., 2017) lead to the formation of a more massive helium core. Additionally, if a hydrogen envelope is still present during the reverse mass transfer, the angular momentum transferred to the primary does not propagate into the core due to the abundance gradient between the core and the hydrogen shell. This results in less angular momentum gain during the reverse mass transfer, and as the primary loses its hydrogen envelope before collapse, it loses the majority of its angular momentum, preventing the formation of an accretion disk. At the same time, the islands of unstable mass transfer from $L_{2}$ overflow and the maximum $\dot{M}$ are expanding.
Moving to $Z=0.2\leavevmode\nobreak\ \mathrm{Z}_{\odot}$, two additional effects are starting to show: short-period systems no longer experience reverse mass transfer, while long-period systems are now experiencing a contact phase, which becomes dominant towards even lower metallicities, as the $Z=10^{-3}\leavevmode\nobreak\ \mathrm{Z}_{\odot}$ panel in Figure 4 shows. In wider binaries, the mass transfer no longer sufficiently strips the primary star for it to reach the Wolf-Rayet wind regime, instead, it finds a thermal equilibrium close to or at its Roche lobe (Klencki et al., 2022; Ercolino et al., 2024). As the companion evolves off its main sequence and also fills its Roche lobe, the system goes into contact, either because the primary is already filling its Roche lobe or because its equilibrium state is near the Roche lobe, such that the radial expansion from mass accretion causes the primary to fill its Roche lobe. On the other hand, short-period binaries no longer experience reverse mass transfer, because the companion remains compact and hydrogen-burning until the primary reaches carbon exhaustion. Due to lower stellar wind mass loss, these accretors retain the gained angular momentum from the initial mass transfer phase. At lower metallicities, the radial expansion of the stars is already less compared to solar metallicity due to the lower opacity, and the increased rotational mixing further limits the radial expansion, which restricts a reverse-mass-transfer phase. The combination of these effects creates a LGRB formation channel that becomes more efficient as metallicity increases, while completely disappearing at metallicities below $0.2\leavevmode\nobreak\ \mathrm{Z}_{\odot}$.
@@ 3.2 Stable reverse mass transfer LGRB rate
Given that most observed LGRBs occur in low-metallicity environments, this indicates that some effect must be limiting a large fraction of our stable reverse mass-transfer model progenitors from producing a LGRB. Theoretically, one would expect that such an effect follows from the required energy for the jet to propagate through the stellar envelope and successfully break out of it (e.g. Matzner, 2003; Bromberg et al., 2011; Salafia et al., 2020). The binding energy of the stellar structure at core carbon exhaustion is less than the values of $E_{\mathrm{BZ}}/\eta_{\phi}$, when $\eta_{\phi}=1$. While the actual energy breakout in general depends on the properties of the stellar envelope and also on those of the jet at launch. Here, for simplicity, we model both as a universal threshold on the jet energy. Because we left the efficiency factor $\eta_{\phi}$ unspecified (see Section 2), we set an energy limit for a successful LGRB of $10^{51}\leavevmode\nobreak\ \mathrm{erg}$ based on the typical breakout energy from jet propagation simulation (for example, see Harrison et al., 2018). In what follows, we explore five values for the efficiency, $\eta_{\phi}$, namely $1$, $0.1$, $10^{-2}$, $10^{-3}$, and $5\times 10^{-4}$.
The total rate of disk formation through the stable reverse-mass-transfer channel is comparable to the local rate density of observed LGRBs (Ghirlanda & Salvaterra, 2022). Given that the inferred fraction of high-metallicity host galaxies is 10%-20%, this requires an efficiency that removes 80%-90% of events. The most restrictive efficiency with $\eta_{\phi}=5\times 10^{-4}$ results in a volumetric rate of ${\simeq}2.01$ Gpc-3 yr-1 at $z=0$; with a total of ${\simeq}15.0$ Gpc-3 yr-1 events below $z<2.5$. With $\eta_{\phi}=10^{-3}$, the reverse-mass-transfer channel rate is around 18% of the latest estimate of the observed cosmic LGRB rate. This is remarkably close to the observational fraction of solar and super-solar host galaxies of 10-20 per cent. However, because the normalisation of the observed rate is uncertain due to selection effects, such as the intrinsic energy of the jet, one should be careful to interpret a comparison of our theoretical rates against these observations. These cannot be constrained without considering all formation channels concurrently, which is beyond the scope of this work. The intrinsic rate density evolution with redshift for different choices in the efficiency is shown in Figure 5. We report their rate densities at $z=0$ and averaged within $z=2.5$ in Table 1.
@@
Figure 5: The redshift evolution of the stable reverse-mass-transfer LGRB event rate density for the IllustrisTNG metallicity and star formation history. The total potential population is indicated in black, while the dashed red line shows the intrinsic LGRB rate from Ghirlanda & Salvaterra (2022) with the best-fit parameters based on observations. The red shaded area is 10% to 20% of this rate. The five green shaded lines indicate the rate density for an energy limit of $10^{51}\leavevmode\nobreak\ \mathrm{erg}$ and different $\eta_{\phi}$.
@@ 4 Discussion
While LGRBs have been found in high-metallicity environments, no collapsar formation channel with a positive metallicity relationship has previously been proposed. We have introduced a new formation channel capable of forming sufficient LGRBs to match observational or theoretical fractional rates in high-metallicity environments. However, much uncertainty remains in this fraction, as well as in the absolute rate of the total LGRBs and this channel. Since the main contributors to this uncertainty are the observational biases and hidden LGRB fraction, which would require a population study that includes all possible LGRB formation pathways in order to be explored, we instead focus on the physics which might impact the presence and occurrence of the reverse-mass-transfer LGRB formation channel.
@@ 4.1 Implications of energy limit
We have implemented a universal energy limit as a proxy for successful jet breakout with different efficiencies, which is equivalent to different energy limits with a constant efficiency, i.e. a $10^{51}\leavevmode\nobreak\ \mathrm{erg}$ limit with a $\eta_{\phi}=10^{-3}$ selects the same progenitors as a $10^{54}\leavevmode\nobreak\ \mathrm{erg}$ limit with a $\eta_{\phi}=1$. For the stable reverse-mass-transfer channel to produce 10-20% of the local LGRB rate, a low disk-to-jet-energy conversion or a high jet-breakout-energy threshold is required. This implies that successful jets are a minority, while choked jets occur an order of magnitude more frequently in the proposed channel. This fraction aligns with Bromberg et al. (2012) and from observed low-luminosity GRBs the inferred intrinsic rate can be up to an order of magnitude higher than their high-luminosity counterpart (Liang et al., 2007; Guetta & Valle, 2007; Virgili et al., 2009). It suggests the existence of a large population of relativistic supernovae with potentially low-luminosity gamma-ray emission caused by these choked jets (see, for example, Soderberg et al., 2006; Wang et al., 2007; Nakar & Sari, 2012).
@@ 4.2 Type Ic broad-line and potential LGRB companions
LGRBs and SN Ic-BL have been associated with one another, but this relationship is not universal; not all Type Ic-BL are accompanied by a LGRB and vice versa (Modjaz et al., 2016; Japelj et al., 2018). Type Ic-BL without a LGRBs have distinct spectral features from those with a LGRB (Modjaz et al., 2016), and could indicate a failed jet breakout (see Corsi & Lazzati, 2021, and references therein). In general, Type Ic-BL are associated more with super-solar host galaxies compared to LGRBs (Japelj et al., 2018), which aligns with the positive metallicity dependence of the stable reverse-mass-transfer LGRB channel. Additionally, if this channel produces the observed high-metallicity population of LGRBs, it produces an order of magnitude more choked jets compared to successful jets (see Section 4.1). Although the observed relative rate of SN Ic-BL to core-collapse SNe is poorly constrained, based on a single SN Ic-BL observation in the volume-limited Lick Observatory Supernova Search, SN Ic-BL make up 4% of stripped-envelope supernovae (Shivvers et al., 2017). Combined with the stripped-envelope supernova rate from Frohmaier et al. (2021), this results in a rough SN Ic-BL rate estimate of ${\sim}300$ Gpc-3 yr-1. Assuming that all disk-forming systems produce a jet and $\eta_{\phi}=10^{-3}$, ${\simeq}86$ choked jets occur per Gpc3 per year at $z=0$ from the stable reverse-mass-transfer channel. This is comparable to the observationally inferred rate and could imply that either more dominant LGRB formation channels must have a smaller ratio of choked-to-successful jets to not overproduce the SN Ic-BL or not all failed LGRB accretion disk systems produce a SN Ic-BL.
@@ 4.3 Jet physics
The relationship between the progenitor star and jet energy remains highly uncertain and general relativistic, radiation, magnetohydrodynamic simulations have shown that the efficiencies depend on the state of the disk (Tchekhovskoy et al., 2010). A popular approach is the magnetically arrested disk (MAD) state because it produces powerful jets even for slowly rotating BHs and causes rapidly spinning BHs to spin down (Jacquemin-Ide et al., 2024). In this state, the mass accretion rate and spin-down timescale determine the jet power. In such disks $\eta_{\phi}=1$, while $\eta_{a}$ can be fitted following Lowell et al. (2024). This can result in efficiencies above 1 due to the extraction of angular momentum from the BH and results in high-energy jets. These even occur for slowly spinning BHs as long as the progenitor has sufficient angular momentum to form an accretion disk (Fryer et al., 2025). However, recent simulations of collapsars show that the accretion rate remains sufficiently high that neutrino cooling keeps the disk thin and the MAD state is not reached when assuming a proto-neutron star is the origin of the BH magnetic field (Gottlieb et al., 2024).
@@
Besides the potential energy budget to form a jet, an essential component in LGRB formation in the collapsar scenario is the jet breakout through the stellar structure. The jet breakout depends on the injected energy and active duration of the central engine, and the density profile of the stellar structure. While the internal stellar density profile at core-carbon exhaustion is available, we do not evolve the final evolutionary phases due to the required computational time, and, thus, do not have a representative profile at collapse available. We verify that the free-fall timescales of the core-carbon exhausted models are similar to observed $T_{90}$ values (See Appendix C). Furthermore, the injected energy depends on several poorly constrained parameters, such as the injection opening angle, the disk mass to jet energy conversion, and the total jet energy emitted in the gamma-ray. A full end-to-end pipeline with computationally intensive hydrodynamical simulations is required to accurately model the link between the progenitor model and the jet properties from the collapsar scenario (Gottlieb et al., 2024).
@@ 4.4 Star formation history
Although the total star formation rate density in the local Universe is relatively well constrained, the metallicity evolution and spread are less understood. It is the high-metallicity-specific star formation rate that gives rise to uncertainty in the shape of the stable reverse-mass-transfer LGRBs and in the weighting of each model over redshift. The former can be seen in Figure 7, where we compare the LGRB rates over redshift from the IllustrisTNG simulation with the two empirical star formation history prescriptions described in Section 2. The metallicity distribution redistributes the weight between populations of different metallicities but does not alter the shape of the potential LGRB energy distribution per metallicity (see Appendix B). The ln-normal metallicity distribution from Neijssel et al. (2019) produces a very peaked metallicity-specific star formation rate, reducing the contribution from the reverse-mass-transfer LGRB channel at high redshift to the LGRB rate, while the higher metallicity and wider distribution from Bavera et al. (2020) allow for more near-solar metallicity star formation at higher redshift.
@@
A stronger effect in the local Universe comes from the energy cuts. Due to the reweighing of each LGRB progenitor, the cumulative potential energy distribution is significantly different per star formation history, thus, the energy cut can suddenly remove more events. For example, the most restrictive cut in Figure 6 reduces the stable reverse-mass-transfer LGRB rate to 2.01 LGRB Gpc-3 yr-1, while the Neijssel et al. (2019) prescription in Figure 16 reduces it to 0.17 LGRB Gpc-3 yr-1. This is due to the energy distribution at each metallicity, which each SFH rebalances, as can be seen in Appendix B. As a result, the choice in mean metallicity and metallicity spread has a major impact on the final rate and the limit of the jet energy.
@@ 4.5 Additional caveats
After the stable reverse-mass-transfer phase, the surface hydrogen fraction of the primary can return to above 0.4, which is the boundary for the Wolf-Rayet winds. As such, the stellar winds for OB stars are applied to this phase of the evolution, which will transition back into the Wolf-Rayet winds as the small hydrogen layer is stripped and the surface hydrogen fraction reaches below 0.4. The transition to the Vink et al. (2001) winds means a lower mass loss in comparison to the Nugis & Lamers (2000) Wolf-Rayet winds are applied. If the mass loss during this phase were treated as a Wolf-Rayet wind, more mass and angular momentum would be lost in the final phase, which would shrink the number of primaries with sufficient angular momentum to produce an accretion disk. On the other hand, the total mass and angular momentum loss strongly depend on how far the core helium burning has progressed when the stable reverse mass transfer takes place. Generally, for potential LGRB progenitors in the stable reverse-mass-transfer channel, the interaction occurs late in the core helium-burning phase of the primary. Otherwise, the hydrogen envelope is not yet fully stripped, which inhibits the transport of angular momentum into the helium core. With no clear observations of hydrogen-accreted helium stars, the exact treatment of their mass loss is unclear.
The first mass transfer in the binary can result in rejuvenation of the companion, where the efficiency of semi-convection affects the parameter range for which this occurs (e.g. Braun & Langer, 1995). The POSYDON v2 grids use $\alpha_{\mathrm{sc}}=0.1$ (for more details, see Andrews et al. (2025)), a relatively inefficient semi-convection, which limits the parameter space for rejuvenation. Furthermore, the stable reverse-mass-transfer channel predominantly occurs around $q\approx 0.95$ at $\mathrm{Z}_{\odot}$. This implies that the companion is far along the main-sequence when the first mass transfer takes place. This limits the effect of the accreted material on the further evolution and the timing of the reverse mass transfer phase. It is unlikely that more efficient semi-convection can lead to more rejuvenation and less stable reverse mass transfer in this regime. At low metallicities, the first mass transfer does affect the companions, and they remain compact until the primary reaches carbon-exhaustion. More efficient semi-convection could result in more rejuvenation, while less efficient semi-convection could limit the rejuvenation and allow for a reverse mass transfer before the primary collapses. However, these interactions are likely to be unstable, since the primary will not be fully stripped at the moment of the interaction. As such, the choice of semi-convection parameter should not affect the overall results of this work.
@@
While the majority of LGRB progenitors from the stable reverse-mass-transfer channel are well into the BH formation regime, their parameter space at low masses is limited by neutron star formation, as shown in the right grid slice in Figure 3. Allowing for more BH formation would increase the contribution of this formation channel. However, the additional parameter space is limited, because the timescale difference between the binary components is too large at low masses for reverse mass transfer to occur. Additionally, the Patton & Sukhbold (2020) prescription uses the average central carbon abundance and carbon-oxygen core mass at helium depletion to determine the explodability. Late-stage binary interactions, such as mass transfer after core helium ignition, can alter the stellar structure sufficiently to boost explodability (Laplace et al., 2021). Despite that, the POSYDON MESA models are evolved until carbon depletion, due to the Patton & Sukhbold (2020) prescription, late-stage mass transfer effects are not considered.
Switching the remnant mass prescription to the rapid or delayed prescriptions from Fryer et al. (2012) should give an indicator of how different prescriptions for estimating the remnant masses affect our results. These prescriptions use the core properties at carbon depletion to determine BH formation through fallback onto a proto-compact object. Figure 8 shows minimal differences between their metallicity bias functions, which propagate to nearly the same cosmic rates. Slightly less stable reverse-mass-transfer LGRBs occur with the Fryer et al. (2012) prescriptions. The star formation history, metallicity distributions, and energy cut have a significantly stronger impact on this rate than what stars collapse as BHs.